OpenStudy (anonymous):

The perimeter of a rectangular garden is 342 feet. If the width of the garden is 78 feet, what is its length?

OpenStudy (angela210793):

Sai-f :P

OpenStudy (saifoo.khan):

Perimeter = 2L + 2B 342 = 2L + 2(78) L = 342 - 156 --------- 2 Solve.

OpenStudy (anonymous):

thankyou ill probally be your bigest fan